### v3

# Torsion in Persistent Homology and Neural Networks

Abstract
We explore the role of torsion in hybrid deep learning models that incorporate topological data analysis, focusing on autoencoders. While most TDA tools use field coefficients, this conceals torsional features present in integer homology. We show that torsion can be lost during encoding, altered in the latent space, and in many cases, not reconstructed by standard decoders. Using both synthetic and high-dimensional data, we evaluate torsion sensitivity to perturbations and assess its recoverability across several autoencoder architectures. Our findings reveal key limitations of field-based approaches and underline the need for architectures or loss terms that preserve torsional information for robust data representation.
1 Introduction
Topological data analysis (TDA) has become increasingly relevant across scientific disciplines, from cancer image analysis to theoretical particle physics [Rabadán2020, thomas2022topological]. In parallel, several efficient and accessible TDA libraries have been developed, such as Ripser, GUDHI, and giotto-tda [Bauer2021Ripser, giottotda, gudhi:urm]. With the growing dominance of neural networks in data-driven applications, TDA techniques have also found their way into machine learning pipelines, particularly in unsupervised settings using autoencoders.
Autoencoders have proven effective for nonlinear dimensionality reduction and unsupervised representation learning in a broad application field [Finke2021Autoencoders, Badsha2020, fengbreast, hooliver]. Recent work has sought to enhance their topological sensitivity by incorporating TDA-inspired components — either through loss functions that reward topological similarity between input and output [moor2021topologicalautoencoders, trofimov2023learningtopologypreservingdatarepresentations], or through architectural modifications tailored to preserve topological structure [carrière2020perslayneuralnetworklayer].
A central topological feature that remains largely overlooked in this context is torsion. While torsion is often associated with abstract examples like the Möbius strip or Klein bottle, it also arises in applied contexts, such as tumor growth models [amster2024topological], Section 9.5 and the patch space of natural images [Carlsson2008LocalBehavior]. Furthermore, recent work [obayashi2023field, kahle2018cohenlenstraheuristicstorsionhomology] suggests that torsional features are likely to occur in high-dimensional data — the kind routinely encountered in deep learning applications.
However, torsion is invisible over field coefficients and can only be detected via integer (persistent) homology, as implied by the universal coefficient theorem ([hatcher], Theorem 3A.3 for homology). Since most hybrid approaches in deep learning rely on field coefficients for computational tractability, they inherently neglect torsional information. This omission is typically unacknowledged, and it is often assumed that the choice of coefficient ring has little or no impact on downstream tasks, explainability or interpretability.
In this work, we challenge that assumption. We show — both computationally and mathematically — that torsion is a fragile but significant topological feature whose presence exposes fundamental limitations of current TDA-enhanced autoencoders. Even architectures that explicitly incorporate TDA loss functions are unable to capture or reconstruct torsion reliably.
1.1 Contributions
Torsion remains a rarely addressed phenomenon in both computational and theoretical topology. Yet it plays a key role in dynamical systems theory, where it serves as a distinguishing feature — for instance, separating models of tumor growth from chaotic attractors such as the Rössler system [amster2024topological]. Despite its relevance, torsion is often treated as a theoretical curiosity and is largely absent from practical applications of topological data analysis (TDA), particularly in deep learning.
Autoencoders, as nonlinear tools for representation learning, offer a promising interface with TDA. They enable topological computations in low-dimensional latent spaces, where standard algorithms would otherwise be computationally infeasible. However, the extent to which these models preserve or distort torsional features remains unknown.
This work is the first to systematically explore torsion phenomena in the context of neural networks. We identify several structural and computational limitations and highlight challenges for future hybrid architectures.
From a mathematical perspective, we show:
  • how torsion is typically lost due to dimensionality reduction in the encoder (Section LABEL:sec:dimred),
  • how torsion that is preserved in the latent space leads to changes in persistence diagrams over field coefficients (Section LABEL:sec:torsreplatsp),
  • and which decoder architectures fail to reconstruct torsion — and in which cases reconstruction becomes possible (Section LABEL:sec:lindec).
In particular, our results imply that, even when working over a coefficient field, one cannot expect a trained autoencoder to induce isomorphisms between the homology groups of the input and latent spaces when torsional phenomena are involved. This challenges the assumptions underlying many TDA-inspired loss functions, which often aim to align field-based homological invariants between input, output, or latent representations [moor2021topologicalautoencoders, trofimov2023learningtopologypreservingdatarepresentations].
From a computational perspective, we demonstrate:
  • that torsion is highly unstable under local perturbations of the input data — small changes such as the displacement or removal of a single point can eliminate torsion entirely, and the prime order of torsion classes is sensitive to such modifications,
  • and that torsion reconstruction is not guaranteed, even for architectures explicitly designed to preserve topological features.
To support our claims, we perform experiments on both synthetic and high-dimensional datasets. For illustration, we use the double and triple loop point clouds embedded in three-dimensional Euclidean space ([obayashi2023field], Figure 4(a) and 4(e)), as well as randomly sampled high-dimensional point clouds. Torsion is detected using the algorithm presented in [obayashi2023field], implemented in the homcloud persistent homology software package [homcloud].
1.2 Related Work
Zomorodian, who introduced the so-called standard algorithm for computing persistent homology over an arbitrary coefficient ring in [zomorodian2005computing], was the first to address torsion phenomena from a computational perspective within topological data analysis. This line of inquiry was further developed by Obayashi [obayashi2023field], who examined the impact of torsion phenomena on invariants computed over a coefficient field in persistent homology, both from a computational and a mathematical standpoint.
Computational results also indicate that torsion is a non-negligible phenomenon in high-dimensional data sets. Building on the model developed by Erdős and Rényi for generating one-dimensional random simplicial complexes [erdos1959on, erdos1960evolution], Linial and Meshulam introduced a model for generating higher-dimensional random simplicial complexes [linial2006homological, meshulam2009homological]. As shown by Kahle et al. [kahle2018cohenlenstraheuristicstorsionhomology], there is almost always a moment in this random complex process at which a “burst of torsion” appears, marking a critical topological transition.
2 Persistent Homology
We define persistent homology with regard to its application to machine learning settings. In this context, we understand a data set $\mathcal{D}=\{x_{i}\in\mathbb{R}^{d}\;|\;1\leq i\leq n,\;n\in\mathbb{N}\}$ as a finite (Euclidean) metric space, also called a point cloud. On this data set we construct a finite geometric simplicial complex – a topological space constructed by gluing together dimensional simplices (points, line segments, triangles, tetrahedrons, etc.). We impose a total order
$K:\emptyset=K_{0}\subset K_{1}\subset...\subset K_{N}$ (1)
on the set of simplices called a filtration. In the following, we work predominantly with Vietoris–Rips filtrations (in short: “Rips filtrations”).
Definition 2.1 (Vietoris–Rips Filtration).
Given a point cloud $\mathcal{D}=\{x_{i}\in\mathbb{R}^{d}\;|\;1\leq i\leq n\}$} and positive real numbers $\delta_{0}<\delta_{1}<\cdots<\delta_{m}$ with $\delta_{i}\in\mathbb{R}_{>0}$, the Vietoris–Rips filtration is the nested sequence of simplicial complexes
$\emptyset=V_{\delta_{0}}(\mathcal{D})\subseteq V_{\delta_{1}}(\mathcal{D})\subseteq\cdots\subseteq V_{\delta_{m}}(\mathcal{D})$ (2)
where a $k$-simplex in $V_{\delta_{i}}(\mathcal{D})$ exists if all pairwise distances between its $k+1$ vertices are $\leq\delta_{i}$.
Definition 2.1 introduces the notion of an abstract filtered simplicial complex, a purely combinatorial object that can be endowed with the structure of a topological space via geometric realization ([edelharerbook], Section III.1). For a Vietoris–Rips filtration constructed on a finite metric space $M\subset\mathbb{R}^{d}$, the dimension of its geometric realization may exceed $d$. In the following, we restrict attention to geometric filtered simplicial complexes that are embedded in $\mathbb{R}^{d}$, thereby excluding phenomena such as Vietoris–Rips bubbles [vrplanarpoint]. Our objects of study are thus finite geometric filtered simplicial complexes $K$ constructed from finite metric spaces $M\subset\mathbb{R}^{d}$. The $p$-th persistent homology of a finite-dimensional filtered simplicial complex with coefficients in a commutative ring $R$ is given by
$PH_{p}(K;R):H_{p}(K_{0};R)\rightarrow...\rightarrow H_{p}(K_{N};R),$ (3)
where $H_{p}(K_{i},R)$ denotes the $p$-th simplicial homology group for a subcomplex $K_{i}$, $0\leq i\leq N$, in the filtration. This sequence captures how topological features such as connected components, loops, and voids appear and disappear across different scales. The image of the map $H_{p}(K_{i};R)\to H_{p}(K_{j};R)$ is referred to as a persistent homology group and identifies features that persist across the filtration interval $[i,j)$.
Over $R=\mathbb{Z}$, the simplicial homology groups $H_{p}(K,\mathbb{Z})$ decompose, by the structure theorem for finitely generated abelian groups [dummit2003abstract, gallian2021abstract], into a direct sum of a free part and a finite torsion subgroup.
2.1 Field Coefficients and Persistence Diagrams
For field coefficients, the torsion subgroup vanishes, and persistent homology is fully determined by the ranks of the free parts of the persistent homology groups (field Betti numbers). As a result, its topological invariants admit a simple representation: namely, the persistence intervals $[i,j)$ over which a feature exists within the filtration encode all relevant information. These intervals can be visualized either as points in a persistence diagram—where the $x$-axis denotes the birth time of a feature and the $y$-axis its death time—or as horizontal bars in a so-called barcode diagram. In the following, we denote by $\mathrm{dgm}_{p}(X;R)$ the persistence diagram in homological dimension $p$, with coefficients in the ring $R$, of a topological space $X$ obtained as the topological realization of a filtration of a finite metric space.
In computational data analysis, persistent homology is highly popular because it is stable under noise and sufficiently small perturbations of the data. In other words, the metric distance (bottleneck distance) between the persistence diagrams of two topological spaces that have a small Hausdorff distance is also small [skraba2023wassersteinstabilitypersistencediagrams]. On the other hand, a common criticism for persistent homology is that individual outliers can have significant effects — they can even be considered “deadly” [chazal2014robusttopologicalinferencedistance] in the sense that they significantly change the field Betti numbers of the underlying topological space.
2.2 Torsion
Torsion coefficients annihilate specific elements of the homology group via multiplication. Specifically, for each element $z\in H_{p}(K;\mathbb{Z})$ of the torsion subgroup, there exists an integer $q^{k}$, for $q$ prime, such that
$q^{k}z=0,$
where $q^{k}$ is a torsion coefficient and $q$ its prime divisor. We say that the integral homology group of $K$ for homology dimension $p$ exhibits q-torsion [hatcher].
Remark 2.1.
In a filtered simplicial complex, torsion may occur either in the absolute homology group of a subcomplex $K_{i}$, or in a relative homology group $H_{*}(K_{i},K_{j})$, where $K_{j}\subseteq K_{i}$ are subcomplexes appearing in the filtration of $K$. An instance where torsion arises exclusively in relative homology, but not in any absolute homology group of the filtration, is given in [obayashi2023field], Example 1.5.
For our subsequent discussion, we need the following result from [obayashi2023field].
Theorem 2.2 ([obayashi2023field], Theorem 1.6).
Let $X$ be a topological space obtained as the topological realization of a filtration of a finite metric space, $k$ a coefficient field and $p$ a homology dimension. The persistence diagram $\mathrm{dgm}_{p}(X;k)$ is independent of the choice of coefficient field, if the relative integral homology group $H_{p}(X_{i},X_{j};\mathbb{Z})$ is free for any indices $0\leq i<j\leq N$ in the filtration of $X$ and $H_{p-1}(X_{i};\mathbb{Z})$ is free for any $0\leq i\leq N$.
3 Neural Networks
Neural networks represent a class of machine learning models composed of interconnected computational units, inspired by the architecture of biological neurons. These units, or neurons, are typically organized into layers. A network with more than three layers is called deep. We consider a neural network $f$ as a composition of $N\in\mathbb{N}$ sequentially connected layers $f_{j}$, $1\leq j\leq N$:
$f(x_{i,nn})=f_{N}\circ f_{N-1}\circ\dots\circ f_{2}\circ f_{1}(x_{i,nn}).$ (4)
The composition $f_{N}\circ f_{N-1}\circ\dots\circ f_{1}$ represents the forward propagation of an input $x_{i,nn}$ through the network’s layers, yielding the final output. Thus such a network is also known as a feedforward network.
Let $\mathcal{D}=\{x_{i}\in\mathbb{R}^{d}\;|\;1\leq i\leq n,\;n\in\mathbb{N}\}$ be a data set serving as the input to the neural network. The output $o_{ijk}$ of neuron $k$ in layer $j$ for an input $x_{i}\in\mathbb{R}^{v}$ is computed as:
$o_{ijk}=\sigma\left(W_{jk}\cdot x_{i}+b_{jk}\right),$ (5)
where $\sigma:\mathbb{R}\to\mathbb{R}$ denotes a typically nonlinear activation function. The term $W_{jk}\in\mathbb{R}^{v}$ represents the weight vector associated with neuron $k$ in layer $j$, and $x_{i}\in\mathbb{R}^{v}$ is the input to this layer. Finally, $b_{jk}\in\mathbb{R}$ denotes the bias term associated with neuron $k$ in layer $j$.
Assumption 3.1 (Euclidicity).
For a given layer $f_{j}$ containing $M\in\mathbb{N}$ neurons, the output for an input $x_{i,l}\in\mathbb{R}^{v}$ to the layer is assumed to be a vector $f_{j}(x_{i,l})\in\mathbb{R}^{M}$, defined as:
$f_{j}(x_{i,l})=[o_{ij1},o_{ij2},...,o_{ijM}],$ (6)
where each $o_{ijk}\in\mathbb{R}$ for $1\leq k\leq M$ denotes the output of the $k$-th neuron in layer $j$ for the input $x_{i,l}$.
We denote the term in Eq. (6) as the layer state.
The weights and biases are parameters of the neuronal network that are optimized during an iterative process called training. The difference (also called loss) between an output $y_{i}=f(x_{i,nn})$ associated with a specific input $x_{i,nn}$ to a neuronal network is measured with respect to a pre-selected metric distance (the error metric or loss term). This is selected in a way that a minimum loss would indicate the network to have learned a desirable characteristic of the input data. By default, the mean squared error (MSE) is selected in applications, defined as:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
where $y_{i}$ are the true outputs, $\hat{y}_{i}$ are the predicted outputs and $n$ denotes the cardinality of samples in the input data.
4 Autoencoders
Autoencoders [rumelhart] constitute a network architecture initially introduced for dimensionality reduction [GoodBengCour16]. Via the use of nonlinear activation functions, autoencoders represent a nonlinear dimensionality reduction method.
Definition 4.1 (Autoencoder).
Let $\mathcal{D}\subset\mathbb{R}^{d}$ be an input data set. Under Assumption 3.1, an autoencoder is a neural network comprising two parameterized functions (we assume each given in the form of Eq. (4)), namely, the encoder
$f_{\text{enc}}(\cdot;\theta_{\text{enc}}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{s}$ (7)
and the decoder
$f_{\text{dec}}(\cdot;\theta_{\text{dec}}):\mathbb{R}^{s}\rightarrow\mathbb{R}^{d},$ (8)
where $\theta_{\text{enc}}$ denote the encoder’s parameters and $\theta_{\text{dec}}$ denote the decoder’s parameters respectively. Together they define the autoencoder as a function
$\begin{split}f_{\text{auto}}&:\mathbb{R}^{d}\to\mathbb{R}^{d}\\ f_{\text{auto}}&:=f_{\text{dec}}\circ f_{\text{enc}}.\end{split}$ (9)
The output of the encoder is called the latent space.
The defining characteristic of an autoencoder is that the latent layer contains significantly fewer neurons than the input and output layers, i.e. $s<<d$ in Definition 4.1.
If the encoder does not exhibit stochasticity, the encoder is surjective. However, autoencoders are not injective and thus not bijective if not explicitly constructed so.
With regard to their broad applicability and the need for non-model driven ways for data analysis, there exists a wide variation of autoencoder architectures. We call an autoencoder trained using the MSE loss vanilla. In Appendix LABEL:app:AE, we present two modifications of the loss term inspired from TDA, that we use for subsequent experiments in this work.
5 Torsion and Stability
In an autoencoder, the data undergoes (non-linear) transformations through the weight matrices and activation functions. This raises two questions:
  1. I.
    How do these transformations affect the torsion subgroups in the input data?
  2. II.
    Whether and how does the encoder remove torsion from the data so that it no longer exists in the latent space?
Two factors are crucial for answering these questions: (1) the dimensionality reduction and (2) the shift of the data. Specifically, based on Assumption 3.1, the transformation induced by the weight matrix and the activation function is understood as a shift of the data points in Euclidean space.
5.1 Data Shift
Torsion is sensitive to small changes and perturbations in the underlying data. This becomes apparent already in the following minimal example using a Möbius strip.
Example 5.1 (Möbius Strip).
Figure LABEL:fig:moebiusstrip exhibits an illustration of the minimal triangulation of the Möbius strip. Clearly, the removal of $a_{3}$ results in the removal of the non-orientability and thus the relative torsion in the Möbius strip structure, despite still exhibiting an $H_{1}$–cycle due to the remaining connection at $a_{0}$.